OpenStudy (anonymous):

A circle can be defined by an equation x2 + y2 + ax + by + c = 0. Find an equation of a circle that passes through (-4, 1), (3, 0) and (5,4).

OpenStudy (anonymous):

what's the radius of that circle?

OpenStudy (anonymous):

this is all they gave me

OpenStudy (anonymous):

(-4-a)^2 +(1-b)^2 =r^2 (3-a)^2 +(0-b)^2 =r^2 (5-a)^2 +(4-b)^2 =r^2

OpenStudy (anonymous):

\[x ^{2}+y^{2}-8y-9=0\]